# RATE: Score Reward Models with Imperfect Rewrites of Rewrites

### v1

@@ 8 Conclusion
We rely on reward models to align LLMs to human values, but reward models are black boxes and it is unclear what aspects of the text they are actually rewarding. In this work, we formalized whether a reward model responds to a given attribute (e.g., helpfulness, complexity, sensitivity) through the language of causality. Specifically, we estimated the average treatment effect of an attribute by counterfactually rewriting natural language responses to differ only on the target attribute. Although this rewrite process introduces bias, we account for it using rewrites of rewrites, which, in expectation, cancel out off-target changes. This procedure yields RATE: Rewrite-based Attribute Treatment Estimators.
Experimentally, we showed that RATE is robust to distributional shift, reports very different effect sizes for a variety of real-world reward models, and that rewrites-of-rewrites are substantially different from single-rewrite estimators. Our method computes causal effects of individual attributes on reward models without enumerating all off-target attributes and introduces a procedure to find out what attributes reward models are really rewarding.
Reproducibility Statement
To facilitate reproducibility of our RATE method, we have taken the following measures: (1) Our code implementation, including scripts for producing rewrites, estimating treatment effects, and generating plots, is provided at https://github.com/toddnief/RATE. (2) The datasets used in our experiments (IMDB, ELI5, HelpSteer, HH RLHF) are publicly available. (3) In Appendix B, we provide randomly sampled texts, rewrites, and rewrites of rewrites for each dataset/attribute combination, allowing the reader to qualitatively evaluate our rewrites. (4) All reward models evaluated in this study (i.e., FsfairX-LLaMA3-RM-v0.1, NCSOFT/Llama-3-OffsetBias-RM-8B, ArmoRM) are open-source. (5) We report confidence intervals for all main results to ensure statistical reliability, using a normal distribution because of our large sample size. (6) Section 5 includes tips for creating effective rewrite instructions and documents challenges encountered during the rewrite process, aiding in the reproduction of our methodology. (7) For the synthetic experiments, we provide details on how we induced correlations in Appendix B.
@@ Acknowledgements
Appendix A Proofs
See 1
@@ Appendix A Proofs
First, we’ll prove the unbiasedness and consistency of $\hat{\tau}_{\text{ATT}}$ and $\hat{\tau}_{\text{ATU}}$, and then use these results to prove the same for $\hat{\tau}_{\text{ATE}}$. Throughout, we use $\tilde{\xi}$ and $\tilde{\tilde{\xi}}$ to denote i.i.d. samples from the distribution $P_{\xi}$, where the former comes from the first rewrite and the latter from the rewrite of the rewrite.
1. Unbiasedness and Consistency of $\hat{\tau}_{\text{ATT}}$
Fix a prompt $x$ and response $y$ with $w=1$, omitting superscripts for convenience. We calculate:
$R(x,\text{Re}(\text{Re}(x,y,0),1))-R(x,\text{Re}(x,y,0))$
@@ 1. Unbiasedness and Consistency of $\hat{\tau}_{\text{ATT}}$
$\mathbb{E}[R(x,\text{Re}(\text{Re}(x,y,0),1))-R(x,\text{Re}(x,y,0))]$ $=\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R(x,y(1,z,\tilde{\tilde{\xi}}))-R(x,y(0,z,\tilde{\xi}))]$
$=\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R_{W,Z}(x,1,z)+R_{\xi}(x,\tilde{\tilde{\xi}})-R_{W,Z}(x,0,z)-R_{\xi}(x,\tilde{\xi})]$
@@
$\mathbb{E}[\hat{\tau}_{\text{ATT}}]=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=1]=\text{ATT}$
For consistency, by the law of large numbers, as $n_{1}\to\infty$:
@@
2. Unbiasedness and Consistency of $\hat{\tau}_{\text{ATU}}$
Similarly, for $w=0$, we calculate:
$R(x,\text{Re}(x,y,1))-R(x,\text{Re}(\text{Re}(x,y,1),0))$
which has expected value:
$\mathbb{E}[R(x,\text{Re}(x,y,1))-R(x,\text{Re}(\text{Re}(x,y,1),0))]$ $=\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R_{W,Z}(x,1,z)+R_{\xi}(x,\tilde{\xi})-R_{W,Z}(x,0,z)-R_{\xi}(x,\tilde{\tilde{\xi}})]$
$=R_{W,Z}(x,1,z)-R_{W,Z}(x,0,z)$
$=R(x,y(1,z,\xi))-R(x,y(0,z,\xi))$
$=R(x,y(1))-R(x,y(0))$
Therefore, as an average over these quantities, we have:
$\mathbb{E}[\hat{\tau}_{\text{ATU}}]=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=0]=\text{ATU}$
For consistency, by the law of large numbers, as $n_{0}\to\infty$:
@@ 2. Unbiasedness and Consistency of $\hat{\tau}_{\text{ATU}}$
3. Unbiasedness and Consistency of $\hat{\tau}_{\text{ATE}}$
The ATE estimator is a weighted average of the ATT and ATU estimators, where the expected value of these weights corresponds to the proportion of treated and untreated samples in the population. Therefore, by the law of total expectation, the expectation of $\hat{\tau}_{\text{ATE}}$ is:
$\mathbb{E}[\hat{\tau}_{\text{ATE}}]$ $=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=1]\cdot P(W=1)$
$+\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=0]\cdot P(W=0)$
$=\mathbb{E}[R(X,Y(1))-R(X,Y(0))]$
[…]
Thus, $\hat{\tau}_{\text{ATE}}$ is an unbiased estimator of the ATE.
For consistency, note that $\hat{\tau}_{\text{ATE}}$ is a weighted average of $\hat{\tau}_{\text{ATT}}$ and $\hat{\tau}_{\text{ATU}}$. As $n_{0},n_{1}\to\infty$, the weights $\frac{n_{1}}{n_{0}+n_{1}}$ and $\frac{n_{0}}{n_{0}+n_{1}}$ converge to $P(W=1)$ and $P(W=0)$ respectively. Therefore, by Slutsky’s theorem and the consistency of $\hat{\tau}_{\text{ATT}}$ and $\hat{\tau}_{\text{ATU}}$:
[…]
Appendix B Experimental Details
Synthetic Experiments
Our synthetic experiments took data from a real-world dataset (IMDB and HelpSteer) and artificially induced a correlation between the target attribute and the off-target attribute. As both the target and off-target attributes are binary, we can easily control the correlation between them. We group the data into the four possible combinations of the target and off-target attributes (e.g., long positive, short positive, long negative, short negative) and then randomly sample from these groups to create a new dataset. We then evaluate the reward model on this new dataset to see how the correlation affects the estimated treatment effect.
Dataset Long Positive Short Positive Long Negative Short Negative $\mathbf{P(\text{long}\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{positive})}$ $\mathbf{P(\text{long}\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{negative})}$
@@ Appendix B Experimental Details
Dataset Helpful Complex Unhelpful Complex Helpful Simple Unhelpful Simple $\mathbf{P(\text{unhelpful}\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{complex})}$ $\mathbf{P(\text{unhelpful}\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{simple})}$
@@
Example Rewrites
@@
I came here from Digg when the collapse came. Before that day, Digg had a far superior look to it.. … (W = 1) I came here from Digg when it collapsed. Digg had a far superior "Web 2.0" CSS look with rounded but… (W = 0) I came here from Digg when it collapsed, and it was quite a journey transitioning from one platform … (0.13708, 0.11329, 0.10987)
@@
I’m sorry, I’m not sure I understand this. Can you clarify what you mean by “task rabbit”? (W = 0) Certainly! "Task Rabbit" is a service that connects people who need help with various tasks to skill… (W = 1) Certainly! "Task Rabbit" is a service that connects people who need help with various tasks to indiv… (0.07668, 0.10774, 0.09397)
@@
Table 9: HH RLHF, Sentiment
@@
The line, of course, is from the Lord’s Prayer - "Thy Will be done on Earth as it is in Heaven". Swe… (W = 1) The line, of course, is from the Lord’s Prayer - "Thy Will be done on Earth as it is in Heaven". Swe… (W = 0) The line, of course, is from the Lord’s Prayer - "Thy Will be done on Earth as it is in Heaven". Swe… (0.09660, 0.08479, 0.10198)
@@
Here is a list format summary of the top 3 big action steps and top 3 little action steps from the c… (W = 1) - Define a "10" marriage: Create a picture of an ideal marriage based on biblical standards. - Set … (W = 0) - Define a "10" marriage: A "10" marriage is one that aligns with biblical principles, characterized… (0.11781, 0.10532, 0.11470)

### v2

@@ 7 Conclusion and Discussion
We rely on reward models to align LLMs to human values, but reward models are black boxes and it is unclear what aspects of the text they are actually rewarding. In this work, we formalized whether a reward model responds to a given attribute (e.g., helpfulness, complexity, sensitivity) through the language of causality. Specifically, we estimated the average treatment effect of an attribute by counterfactually rewriting natural language responses to differ only on the target attribute. Although this rewrite process introduces bias, we account for it using rewrites of rewrites, which, in expectation, cancel out off-target changes (see Figure 3). This procedure yields RATE: Rewrite-based Attribute Treatment Estimators. Empirically, we find both that RATE is effective at estimating causal effects, and that spurious associations cause substantial bias in naive estimators applied to real problems of interest.
Limitations and Future Directions
RATE can only be applied to measure the effect of attributes where a rewrite can approximate an imperfect counterfactual. However, the quality of counterfactuals has no ground truth, so the effectiveness of the rewrite procedure is ultimately a subjective judgment. In practice, we find this not too difficult to judge by simply looking at the generated responses. Nevertheless, it would be an interesting direction for future work to attempt to more formally validate the rewrite quality, e.g., in the style of Bhattacharjee et al. (2024).
RATE only addresses the effect of the attribute on the reward model, without reference to the downstream task that the reward model will be used for. This is both a strength and a limitation. The strength is that it allows us to understand the reward model in isolation, which is significant since a given reward model may be used for many tasks. The limitation is that it is not completely clear how the causal sensitivity the reward model will translate to, e.g., the behavior of a LLM aligned to this reward. It would be an interesting direction for future work to elucidate this connection.
Impact Statement
Our work is a step in the direction of more interpretable reward models. As reward models are used to align large language models, our work may aid in alignment. In particular, RATE may be used to check whether a particular reward model captures desired preferences and values by simply estimating the causal effect of attributes on rewards.
@@ Acknowledgements
Appendix A Reproducibility Statement
To facilitate reproducibility, we have taken the following measures: (1) Our code implementation, including scripts for producing rewrites, estimating treatment effects, and generating plots, is provided at https://github.com/toddnief/RATE. (2) The datasets used in our experiments (IMDB, ELI5, HelpSteer, HH RLHF) are publicly available. (3) In Section E.3, we provide randomly sampled texts, rewrites, and rewrites of rewrites for each dataset/attribute combination, allowing the reader to qualitatively evaluate our rewrites. (4) All reward models evaluated in this study (i.e., FsfairX-LLaMA3-RM-v0.1, NCSOFT/Llama-3-OffsetBias-RM-8B, ArmoRM) are open-source. (5) We report confidence intervals for all main results to ensure statistical reliability, using a normal distribution because of our large sample size. (6) Section E.1 includes tips for creating effective rewrite instructions and documents challenges encountered during the rewrite process, aiding in the reproduction of our methodology. (7) For the synthetic experiments, we provide details on how we induced correlations in Section C.2.
Appendix B Proofs
[…]
@@ Appendix B Proofs
First, we’ll prove the unbiasedness and $\sqrt{n}$-consistency of $\widehat{\text{ATT}}_{\text{RATE}}$. The argument for $\widehat{\text{ATU}}_{\text{RATE}}$ follows by symmetry. Then, we can use these results to prove the same for $\widehat{\text{ATE}}_{\text{RATE}}$. Throughout, we use $\tilde{\xi}$ and $\tilde{\tilde{\xi}}$ to denote i.i.d. samples from the distribution $P_{\text{Re}}$, where the former comes from the first rewrite and the latter from the rewrite of the rewrite.
1. Unbiasedness and $\sqrt{n}$-Consistency of $\widehat{\text{ATT}}_{\text{RATE}}$ Fix a prompt $x$ and response $y$ with $w=1$, omitting superscripts for convenience. Then by our latent variable model, $y=Y(1,z,v)$ for some realizations $z$ and $v$ of $Z$ and $\xi$. We calculate:
$R(x,\text{Re}(\text{Re}(y,0),1))-R(x,\text{Re}(y,0))$
@@
$\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R(x,y(1,z,\tilde{\tilde{\xi}}))-R(x,y(0,z,\tilde{\xi}))]$ $=\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R_{W,Z}(x,1,z)+R_{\xi}(x,\tilde{\tilde{\xi}})]$
$-\mathbb{E}_{\tilde{\xi},\tilde{\tilde{\xi}}\sim P_{\text{Re}}}[R_{W,Z}(x,0,z)+R_{\xi}(x,\tilde{\xi})]$
@@
$\mathbb{E}[\widehat{\text{ATT}}_{\text{RATE}}]=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=1]=\text{ATT}$
For $\sqrt{n}$-consistency, note that $R(\cdot,\cdot)$ is bounded, so its variance is bounded. As the ${x^{i},y^{i}}$ are i.i.d., so are the $R(x^{i},y^{i})$. Thus, $\widehat{\text{ATT}}_{\text{RATE}}$ is an average over $n_{1}$ i.i.d. random variables with finite variance, implying:
$\sqrt{n_{1}}(\widehat{\text{ATT}}_{\text{RATE}}-\text{ATT})=O_{p}(1)$
@@
$\sqrt{n}(\widehat{\text{ATT}}_{\text{RATE}}-\text{ATT})=O_{p}(1)$
2. Unbiasedness and $\sqrt{n}$-Consistency of $\widehat{\text{ATUx}}_{\text{RATE}}$ By the same argument as for ATT and since $P(W=0)\in(0,1)$:
$\mathbb{E}[\widehat{\text{ATU}}_{\text{RATE}}]=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=0]=\text{ATU}$
and
$\sqrt{n}(\widehat{\text{ATU}}_{\text{RATE}}-\text{ATU})=O_{p}(1)$
3. Unbiasedness and $\sqrt{n}$-Consistency of $\widehat{\text{ATE}}_{\text{RATE}}$ The ATE estimator is a weighted average of the ATT and ATU estimators. By the law of total expectation:
$\mathbb{E}[\widehat{\text{ATE}}_{\text{RATE}}]$ $=\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=1]\cdot P(W=1)$
$+\mathbb{E}[R(X,Y(1))-R(X,Y(0))|W=0]\cdot P(W=0)$
$=\mathbb{E}[R(X,Y(1))-R(X,Y(0))]=\text{ATE}$
For $\sqrt{n}$-consistency, we can write:
$\sqrt{n}(\widehat{\text{ATE}}_{\text{RATE}}-\text{ATE})$ $=\frac{n_{1}}{n}\sqrt{n}(\widehat{\text{ATT}}_{\text{RATE}}-\text{ATT})+\frac{n_{0}}{n}\sqrt{n}(\widehat{\text{ATU}}_{\text{RATE}}-\text{ATU})$
Since:
@@
$\sqrt{n}(\widehat{\text{ATT}}_{\text{RATE}}-\text{ATT})=O_{p}(1),\sqrt{n}(\widehat{\text{ATU}}_{\text{RATE}}-\text{ATU})=O_{p}(1)$
By Slutsky’s theorem:
$\sqrt{n}(\widehat{\text{ATE}}_{\text{RATE}}-\text{ATE})=O_{p}(1)$
Appendix C Additional Semi-Synthetic Experiment Details
C.1 Additional Semi-Synthetic Experiment: HelpSteer
In a third synthetic experiment, we show that the RATE estimator is stable under distributional shift even when the expected treatment effect is non-zero. In this experiment, we are evaluating the effect of helpfulness on the ArmoRM reward model using the HelpSteer dataset (sample size $n=5148$). Each example in HelpSteer includes human ratings on five attributes using a Likert-5 scale: Helpfulness, Correctness, Coherence, Complexity, and Verbosity. (Wang et al., 2023). We expect “helpfulness” to have a positive treatment effect on reward scores.
In this synthetic experiment, we introduce a correlation between “complexity” and “helpfulness” into the evaluation dataset. While we expect the treatment effect of “helpfulness” to be positive, we should see a constant treatment effect even under distributional shift. In Figure 6, we can see that the naive treatment effect increases as we introduce a spurious correlation between complexity and helpfulness into the evaluation data, while the RATE estimators remain much closer to constant. Again, we note that the double-rewrite estimator is more stable than the single-rewrite estimator.
Figure 6: We estimate the treatment effect of a concept on a reward model under distributional shift using three estimators: naive (blue), $\widehat{ATE}(\text{Rewrite})$ (green), and $\widehat{ATE}(\text{Rewrite}^{2})$ (orange). We see that the causal estimators remain near constant while the naive estimator changes as the distribution shifts. 95% confidence intervals are shown. We estimate the effect of “helpfulness” on ArmoRM scores using the HelpSteer dataset under distribution shift. We expect the “ground truth” effect of “helpfulness” to be positive and to remain unchanged as we add a spurious correlation between “helpfulness” and “complexity” in the evaluation data. The naive estimator (blue) shows an increasing effect size as we increase the correlation between “complexity” and “helpfulness” while both the single- and double-rewrite estimators (green, orange) remain near constant.
C.2 Synthetic Experiment Implementation
Our semi-synthetic experiments took data from a real-world dataset (IMDB and HelpSteer) and artificially induced a correlation between the target attribute and the off-target attribute. As both the target and off-target attributes are binary, we can easily control the correlation between them. We group the data into the four possible combinations of the target and off-target attributes (e.g., long positive, short positive, long negative, short negative) and then randomly sample from these groups to create a new dataset. We then evaluate the reward model on this new dataset to see how the correlation affects the estimated treatment effect.
Dataset Long Positive Short Positive Long Negative Short Negative $\mathbf{P(\text{long}\mid\text{positive})}$ $\mathbf{P(\text{long}\mid\text{negative})}$
@@ C.2 Synthetic Experiment Implementation
Dataset Helpful Complex Unhelpful Complex Helpful Simple Unhelpful Simple $\mathbf{P(\text{unhelpful}\mid\text{complex})}$ $\mathbf{P(\text{unhelpful}\mid\text{simple})}$
@@
Appendix D Additional Figures
D.1 Single vs. Double-Rewrite Estimates
There are significant differences between the single-rewrite estimator and the double-rewrite estimator. Each subplot shows the ATE, ATT, and ATU for a different reward model and attribute. Hence, using double rewrites is crucial for estimating the true treatment effect of an attribute on a reward model.
Figure 7: Using RATE (rewrites of rewrites) rather than just rewrites changes the estimated treatment effects. Here we compare treatment effects of complexity, using data from HelpSteer.
Figure 8: Using RATE (rewrites of rewrites) rather than just rewrites changes the estimated treatment effects. Here we compare treatment effects of helpfulness, using data from HelpSteer.
Figure 9: Using RATE (rewrites of rewrites) rather than just rewrites changes the estimated treatment effects. Here we compare treatment effects of length, using data from ELI5 and IMDB.
Figure 10: Using RATE (rewrites of rewrites) rather than just rewrites changes the estimated treatment effects. Here we compare treatment effects of sentiment, using data from IMDB and HH-RLHF.
Appendix E Example Rewrites
E.1 Iterating the Rewrite Instructions
One surprising behavior we encountered is that, when the example response in need of a rewrite was phrased as a question, the LLM would often answer the question rather than rewriting it. Based on this, we included explicit instructions not to answer questions but, rather, to rewrite them for the HH-RLHF dataset (see Table 5).
Attribute W=1 W=0 Rewrite Prompt
Table 5: Example rewrite prompts from experiments with sentiment and length as the target attribute. For the ELI5 dataset, some of the responses were phrased as questions, so we instructed our rewriter (GPT-4o) not to answer the question and instead rewrite it.
E.2 Ill-defined Target Attributes
In some cases, the target attribute is not well-defined. For example, in the HH-RLHF dataset, neutral questions have no clear sentiment, and consequently the rewrites add strange syntax: “annoyingly the same size” and “frustratingly square” (see Table 6). We qualitatively check for these cases, and only use datasets where the target attribute is consistently well-defined.
Prompt Original (W = 0) Rewrite of Rewrite (W = 0)
Original (W = 1) Rewrite (W = 0)
OK ist not hte best film I’ve ever sene but at the same time I’ve been able to sit and watch it TWICE!!! story ilne wsa pretty awful… Not the best film I’ve ever seen, but at the same time I’ve been able to sit and watch it TWICE!!! Storyline was pretty awful …
Table 7: When asked to rewrite a movie review which starts with a vowel ($W=1$) to instead start with a consonant ($W=0$), GPT-4o also fixes typos, an off-target attribute. The original sample (left) is the first sentence of an IMDB review which had 30% of its words replaced with typos (see Section 5.1).
Original Rewrite Rewrite of Rewrite
W = 0, Reward: 0.14 W = 1, Reward: 0.12 W = 0, Reward: 0.16
W = 0, Reward: 0.13 W = 1, Reward: 0.17 W = 0, Reward: 0.16
W = 1, Reward: 0.135 W = 0, Reward: 0.134 W = 1, Reward: 0.139
Table 8: Whether for a rewrite or a rewrite-of-a-rewrite, GPT-4o tends to use well-formatted text and a slightly formal tone. Here, W is length; samples are drawn from the ELI5 dataset, scored using ArmoRM, and truncated to 100 characters for display. The first was selected for illustrative purposes, the latter two were randomly selected from the dataset.
E.3 Excerpts of Samples
@@
I came here from Digg when the collapse came. Before that day, Digg had a far superior look to it.. … (W = 1) I came here from Digg when it collapsed. Digg had a far superior ”Web 2.0” CSS look with rounded but… (W = 0) I came here from Digg when it collapsed, and it was quite a journey transitioning from one platform … (0.13708, 0.11329, 0.10987)
@@
I’m sorry, I’m not sure I understand this. Can you clarify what you mean by “task rabbit”? (W = 0) Certainly! ”Task Rabbit” is a service that connects people who need help with various tasks to skill… (W = 1) Certainly! ”Task Rabbit” is a service that connects people who need help with various tasks to indiv… (0.07668, 0.10774, 0.09397)
@@
Table 10: HH-RLHF, Sentiment
@@
The line, of course, is from the Lord’s Prayer - ”Thy Will be done on Earth as it is in Heaven”. Swe… (W = 1) The line, of course, is from the Lord’s Prayer - ”Thy Will be done on Earth as it is in Heaven”. Swe… (W = 0) The line, of course, is from the Lord’s Prayer - ”Thy Will be done on Earth as it is in Heaven”. Swe… (0.09660, 0.08479, 0.10198)
@@
Here is a list format summary of the top 3 big action steps and top 3 little action steps from the c… (W = 1) - Define a ”10” marriage: Create a picture of an ideal marriage based on biblical standards. - Set … (W = 0) - Define a ”10” marriage: A ”10” marriage is one that aligns with biblical principles, characterized… (0.11781, 0.10532, 0.11470)